9. If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
A. 2 minutes and 44 seconds
B. 2 minutes and 58 seconds
C. 3 minutes and 10 seconds
D. 3 minutes and 26 seconds
E. 4 minutes and 15 seconds
10. If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
A. 0.8 days
B. 1.09 days
C. 1.23 days
D. 1.65 days
E. 1.97 days


I dont know how to do this type of question. Can you PLEASE explain?

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(6693)

9. If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
:
Here's an easy way to do this:
Let t = time (in min) required when they work together to mix 20 drinks
Let the completed job = 1; (the mixing of 20 drinks)
:
Each person does a fraction of the job and their fractions add up to 1
+ + = 1
Multiply equation by 30 to get rid of he denominators, resulting in:
6t + 3t + 2t = 30
:
11t = 30
t = = 2 & 8/11 min,
Change 8/11 min to seconds; 8/11 * 60 = 43.636
t = 2 min 44 sec
:
:
10. If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
:
Same on this one.
Let t = time (in days) working together
Let the completed job = 1
;
+ + = 1
:
The common denominator will be 12, multiply the equation by 12 and see if you can do this one yourself. If you can't, let me know

Answer by Alan3354(6094)

9. If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
A. 2 minutes and 44 seconds
B. 2 minutes and 58 seconds
C. 3 minutes and 10 seconds
D. 3 minutes and 26 seconds
E. 4 minutes and 15 seconds
-------------------------
Find the LCD of the times, which is 30. Figure how many drinks each would make in 30 minutes:
Steve: 20 in 5 is 120 in 30
Sue: 20 in 10 is 60 in 30
Jack: 20 in 15 is 40 in 30
Add up the drinks: 120+60+40 = 220 drinks in 30 minutes
Now divide by 11 to see how long it takes to make 20 drinks.
220/11 --> 30/11 minutes
30/11 = 2.727 minutes = 2' 44"
Answer A
------------------------------------------
10. If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it?
A. 0.8 days
B. 1.09 days
C. 1.23 days
D. 1.65 days
E. 1.97 days
Instead of days per job, use jobs per day.
For Sam, it's 1/4 per day
for Lisa, it's 1/6
for Tom, it's 1/2
Add up the jobs per day they can do: 1/4 + 1/6 + 1/2
= 3/12 + 2/12 + 6/12 = 11/12
Together, they can do 11/12 jobs per day. Inverting that give the days per job, 12/11 days.
That's B, 1.09 days